# Inductive Learning of Logical Theories with LLMs: A Complexity-graded Analysis

### v1

@@ Abstract
This work presents a novel systematic methodology to analyse the capabilities and limitations of Large Language Models (LLMs) with feedback from a formal inference engine, on logic theory induction. The analysis is complexity-graded w.r.t. rule dependency structure, allowing quantification of specific inference challenges on LLM performance. Integrating LLMs with formal methods is a promising frontier in the Natural Language Processing field, as an important avenue for improving model inference control and explainability. In particular, inductive learning over complex sets of facts and rules, poses unique challenges for current autoregressive models, as they lack explicit symbolic grounding. While they can be complemented by formal systems, the properties delivered by LLMs regarding inductive learning, are not well understood and quantified. Empirical results indicate that the largest LLMs can achieve competitive results against a SOTA Inductive Logic Programming (ILP) system baseline, but also that tracking long predicate relationship chains is a more difficult obstacle than theory complexity for the LLMs.
Inductive Learning of Logical Theories with LLMs:
A Complexity-graded Analysis
João Pedro Gandarela2†,  Danilo S. Carvalho3,  André Freitas1,2,3 1 Department of Computer Science, University of Manchester, United Kingdom 2 Idiap Research Institute, Switzerland 3 National Biomarker Centre, CRUK-MI, Univ. of Manchester, United Kingdom {firstname.lastname}[@manchester.ac.uk]|[@idiap.ch]
1 Introduction
The integration of Large Language Models (LLMs) with formal methods stands out as a promising frontier in the field of Natural Language Processing. It is an avenue for improving model inference control and explainability, by both complementing the content flexibility of Large Language Models (LLMs) with the systematicity of symbolic/formal systems Quan et al. (2024a, b) and by using well-defined formal settings to assess the underlying inference properties of the model.
Inductive Logic Programming (ILP) is a subfield of symbolic AI which focuses on methods that can derive (generalise) theories to explain observed facts and rules Muggleton (1991); Nienhuys-Cheng and de Wolf (1997). Addressing inductive learning over complex sets of facts and rules, poses unique challenges for current autoregressive LLMs, as they do not operate over data symbolically, rather combining an extensive set of structural an semantic signals to approximate the most probable answer in a generative fashion.
@@ 1 Introduction
RQ1. To what extent the combination of an LLM with feedback from a formal inference engine can compare to a SOTA inductive logic programming (ILP) system in logic theory induction, w.r.t. inference quality at different degrees of complexity?
RQ2. How does the complexity of the target theories affect inference quality of LLMs for inductive reasoning?
Figure 1: The proposed method to evaluate theory induction with an LLM in Prolog based on background knowledge and training examples. The process starts with a prompt generator $(c)$ that formulates prompts for an LLM $(a)$. Both the background knowledge and training sets are parameterised by different noise and rule complexities levels: Chain, Rooted Directed Graph (DG), Disjunctive Rooted DG, and Mixed. The LLM generates theories, which are then evaluated by a logic program interpreter $(b)$. The evaluation feedback, including accuracy, precision, recall, and F1 scores, as well as wrongly classified examples, is used to refine the prompts iteratively. We analyse and categorise the generated theories according to their complexity $(d)$.
In order to address these RQs, we propose a method for combining iterative theory refinement on stock LLMs (i.e., zero-shot inference), a formal ILP inference engine and a synthetic generator for inductive reasoning datasets, in order to perform systematic evaluations of the LLM induced theories, using state-of-the-art ILP solvers as a baseline. Moreover, to quantify the extent in which LLMs can address ILP tasks, the evaluation is graded wrt. the dependency complexity of the target rulesets. Figure 1 schematises the proposed approach.
@@
    (Empirical) A detailed empirical analysis on the strengths and limitations of SOTA LLMs regarding logic theory induction generation, according to target ruleset complexity.
@@
The remainder of the paper is organised as follows: Section 2 formalises the relevant ILP concepts, dataset generation, and LLM used in the paper. Section 3 describes the proposed method and algorithms. Section 4 presents the experimental procedures, results and discussion, followed by related work (5) and conclusions (6).
2 Inductive Learning, Complexity & Datasets
In this section we introduce the target task and the typology of inductive learning complexity classes, as well as the dataset generation process.
2.1 Inductive Logic Programming
@@
Given: background knowledge (BK), a set of logical clauses that represent prior knowledge about the domain a set of positive examples ($E^{+}$), a set of ground facts (instances) which the learned theory should entail (i.e., these are examples that should be true according to the theory), a set of negative examples ($E^{-}$), a set of ground facts which the learned theory should not entail (i.e., these are examples that should be false according to the theory). Find a hypothesis $H$ (a set of logical clauses) such that:
Completeness: For every example $e\in E^{+}$, $H\cup BK\models e$, meaning the hypothesis $H$, together with the background knowledge $BK$, should entail all positive examples.
Consistency: For every example $e\in E^{-}$, $H\cup BK\not\models e$, meaning the hypothesis $H$, together with the background knowledge $BK$, should not entail any negative examples.
Formally, an ILP system seeks a hypothesis $H$ that satisfies:
@@
The learned hypothesis $H$ should thus be a logical theory that explains the positive examples and excludes the negative examples, based on the given background knowledge.
2.2 Theory complexity
Inductive learning can be organised according to different classes of complexity, which involves a typology of the structural complexity of the problem. Moreover, variables such as the amount and type of noise within the evidence set (such as the number of incorrect or missing facts or completeness levels) can be integrated within this setting. Following the typology of Cornelio and Thost (2021) four base categories of rules can be introduced based on their dependencies: Chain, Rooted Directed Graph (RDG), Disjunctive Rooted DG, and Mixed. Each category represents a hierarchical generalisation, with each step encompassing a broader range of structures. Starting with CHAIN, which represents a linear composition of predicates, RDG is a generalisation of CHAIN, meaning it includes all chain structures but also accommodates more complex dependencies. Moving up, DRDG generalises RDG by incorporating directed relationships, thus offering a more extensive representation of dependencies. Finally, Mixed contains connected components from CHAIN, RDG, and DRDG. Each progression from CHAIN to MIXED represents a step towards greater inclusivity and complexity in the types of structures captured.
The complexity classes and their characteristics are summarised in Table 1. A detailed description of each class is provided in the supplementary material (Appendix A.1).
@@ 2.2 Theory complexity
Table 1: Characteristics for each dataset category. # Parents refers to the number of rules that each rule can deduce relevant facts to. Recursive refers to whether a predicate in the head of a rule can also occur in the body. Alt. rules indicate whether a predicate can be deduced by alternative rules.
2.3 Dataset synthesis
In order to generate datasets for rigorous analysis, this study employed the RuDaS tool Cornelio and Thost (2021) to systematically vary parameters such as noise, open-world degree, and missing data. By adjusting these factors in conjunction with the category parameter, it is possible to ensure comprehensive coverage of different structural configurations and complexity levels.
@@ 2.3 Dataset synthesis
  • The minimum and maximum number of Directed Acyclic Graphs (DAGs) were both set to 1 (mindags = 1, maxdags = 1).
  • Noise levels were systematically varied at intervals of 0.1, 0.2, and 0.3.
  • The percentage of missing data (missing) and open world degree (owa) were similarly varied across 0.1, 0.2, and 0.3.
  • The category parameter was set to cover all the complexity classes described in the previous section, and listed in Table 1. The distribution of each category in the synthesised dataset is an independent hyperparameter, discussed in Section 4.
Further details regarding the dataset generation can be found in Appendix E.
3 Proposed Approach
The proposed approach can be divided in two parts: iterative refinement and graded evaluation. They form a systematic evaluation loop, covering all the complexity classes described in the previous section, for a given set of LLMs and dataset synthesis parameters.
3.1 Iterative Refinement
@@
(a) A language model capable of generating a theory $H$, based on background knowledge, positive and negative examples, and hypothesis search, provided as a prompt text in a logic program language. Typically an LLM with structured (e.g., code) generation capabilities. (b) A logic program interpreter. We use Prolog Warren et al. (2023) as the logic program language. (c) A prompt generation component, that interleaves logical programming language with natural language queries designed to drive the theory induction responses. The logical programming language expresses background knowledge and the relevant outputs of the program interpreter. (d) An evaluation module, that uses the logic program interpreter to execute the generated theory $H$ as logical rules, and computes a set of evaluation metrics.
@@ 3. Evaluation of Hypothesis: Checking for the completeness and consistency conditions:
from which precision (P) ($\frac{TP}{TP+FP}$), recall (R) ($\text{Recall}=\frac{TP}{TP+FN}$) and F1-score (F1) ($2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$) can be generated.
4. Theory refinement: Following an initial evaluation, the LM is tasked to refine the induced theory iteratively. Each refinement round involves adjusting the theory based on feedback from the Prolog interpreter validation. The refinement aims to improve the theory’s performance by addressing misclassifications and enhancing its predictive capabilities. If $H$ does not satisfy completeness and consistency, update the input representation based on feedback and generate a new hypothesis using the language model, given the Feedback Context $\leftarrow\{FP,FN,P,R,F1\}$ and the final prompt input Input $\leftarrow$ (Refinement Prompt + Context + Feedback Context): $H\leftarrow\mathcal{LM}(\text{Input})$.
@@
3.2 Graded evaluation
A synthetic data generator is used to control for the input parameters of the ruleset complexity, namely: categorical distribution (CHAIN, RDG, DRDG, etc.), background knowledge, positive examples and negative examples as well as the amount of noise introduced within the dataset.
1. Categorised Learning Sets: Consisting of $C$: set of ruleset complexity categories (e.g., CHAIN, RDG, DRDG, etc.), $N$: set of noise levels, $S$: number of samples per combination of $C$ and $N$. For each $c\in C$ and $n\in N$, generate $S$ datasets $\{D_{c,n,i}\mid i=1,\ldots,S\}$ where each dataset $D_{c,n,i}$ includes:
@@ 3.2 Graded evaluation
4 Experiments
In order to answer the research questions, a set of experiments was elaborated to systematically analyse the theory inducing capabilities of a set of most popular open-source LLMs and two versions of GPT, with the proposed approach, having the state-of-the-art ILP system Popper Cropper and Morel (2021) as a baseline. The tests covered all data categories discussed on Section 2, allowing a graded analysis w.r.t. the expected level of induction complexity and tolerated noise.
4.1 Experimental Setup & Dataset
For each data category, five datasets were generated using RuDaS Cornelio and Thost (2021). The size of each dataset was set to $XS$ ($min=50$, $max=100$, $support=3$), and noise, missing, and open world were all set to 0.1, then all set to 0.2, and finally all set to 0.3. This resulted in 105 datasets in total, with 35 datasets for each rate. Subsequently, two methods are used to induce a theory for each dataset: (1) employing Popper, with NuWLS Chu et al. (2023) and WMaxCDCL, varying its time limit parameter from 10 to 800 seconds; (2) applying the proposed iterative LM theory refinement method (Section 3), with parameters $Max_{iter}=4$ and $MT_{thresh}=1.0$. Three different LLM models were used for (2): Open AI [^1] https://openai.com/’s model GPT-4o [^2] https://openai.com/index/hello-gpt-4o, Mistral AI [^3] https://mistral.ai/’s Mixtral-8x7B [^4] https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2023), and Google’s Gemma [^5] https://huggingface.co/google/gemma-7b-it Team et al. (2023).
@@ 4.1 Experimental Setup & Dataset
Table 2: Statistics for each dataset category. A detailed description of each can be found in Section 2.
@@
The experiment used the OpenAI service for GPT models. For Popper, Llama3-8B-Instruct, Gemma-7B-It and Mixtral-8x7B-Instruct-v0.1, it was conducted on a computer with an Intel(R) Xeon(R) Gold 5217 CPU @ 3.00GHz, 188GB RAM, and 2x NVIDIA RTX A6000 (48GB VRAM) GPUs. The software used was CUDA 12.3, PyTorch 2.2.2, and Transformers 4.41.2. Prompt templates used were included in the supplementary material (Appendix C).
4.2 Results & Discussion
Overall results for F1 are presented in Figure 2. We additionally report on processing times as a measure of practical interest in Figure 3. We present the values obtained in detail in the supplementary material (Appendix B, in Tables 4 and 5). Gemma-7B-It results are not included as it failed to generate valid theories. The results reveal significant insights into LLM capabilities and limitations w.r.t. theory induction, which are summarised as follows:
Figure 2: F1 score trends across categories. Different models (GPT-4o, Llama3 8b instruct, Popper, and Mixtral-8x7B-Instruct-v0.1) under varying noise levels and categories reveal distinct performance patterns. GPT-4o demonstrates stable accuracy yet sensitivity to noise, particularly in complex rule-based categories like RDG and DRDG. Mixtral-8x7B-Instruct-v0.1 exhibits mixed results with notable variability across categories particularly in more complex tasks. Llama3 8b instruct struggles with low scores, indicating challenges in reasoning and theory generation.
@@ 4.2 Results & Discussion
Inducing theories on long relation chains is the major obstacle for LLMs, rather than dependency complexity.
With the CHAIN category being the least complex and one of the most consistently solved by Popper, but none of the tested LLMs was able to overcome the baseline performance at all noise levels (Figure 2 CHAIN category). This suggests that such models have a limited capacity of tracking long relationship chains of independent predicates. A part of RQ2 answer.
@@
Upon increasing the iteration limits from 1 to 4, it was found that the metrics can either increase or decrease non-monotonically. Thus $Max_{iter}$ was set to 4 and the iteration with the best accuracy is taken as the final result.
Performance is remarkably lower on complex rule sets at moderate noise levels.
Responses for complex categories, such as RDG and DRDG display higher variance and LLM hallucination artifacts, such as valid rules containing predicates that do not exist in the rule set. We present an error analysis in Section 4.3. For instance, a comparison of the results for the RDG category generated by GPT-4o under noise levels set to 0.1 and 0.3 reveals a significant decline in performance, with the F1-score dropping from 0.75 to 0.22. A comparable pattern is observed with GPT-3.5 Turbo for RDG and DRDG and Mixtral RDG in the presence of elevated noise levels, with GPT-3.5 Turbo scores going from 0.56 to 0.0, 0.28 to 0.08, and Mixtral-8x7B going from 0.60 to 0.0.This complements the answer to RQ2. Further details are included in the supplementary material (Appendix B).
@@
While Popper is a serial algorithm, the parallel nature of transformer-based LLMs allows them to operate at times about 3 orders of magnitude lower, given enough computational resources. This can be observed in Figure 3, for all complexity classes and all tested noise intervals.
4.3 Error analysis
@@
[fontsize=]prolog theory :- p(X, Y), pos(p0(X, Y)) - positive. p(X, Y), neg(p0(X, Y)) - negative. \+ p(X, Y), pos(p0(X, Y)) - false. \+ p(X, Y), neg(p0(X, Y)) - true.
@@
[fontsize=]prolog theory :- p(X, Y) :- p1(X, Y); p3(X, Y); p4(X, Y); p7(X, Y); p8(X, Y); p0(X, Y), not neg(p(X, Y)), (pos(p(X, Y)) - true; fail).
The response contains the head of the clause "theory," as well as the predicates "p" and "pos", which do not exist in the BK. Table 3 presents a distribution of error categories for the analysed models. A more detailed analysis of the models outputs is included in the supplementary material (Appendix G).
@@
5 Related Work
Neural symbolic computation combines neural networks with symbolic methods to enhance AI reasoning capabilities. Yang et al. (2017) introduced Neural Logic Programming, An end-to-end differentiable model integrating neural networks with logic programming. Within the LLM-Symbolic space, Wan et al. (2024) developed LogicAsker, which evaluates and improves LLMs’ logical reasoning using propositional and predicate logic. It identifies reasoning failures and enhances capabilities through in-context learning. Within the context of symbolic toolformers over LLMs, Quan et al. (2024a, b) proposed methods of improving explanatory reasoning with the support of formal iterative cycles using both logical solvers and theorem provers for supporting more controlled step-wise reasoning.
Despite these advancements at the interface of LLM-based reasoning and formal controls, it is unclear the extent and the conditions in which LLMs can perform formal reasoning Huang and Chang (2023). Sinha et al. (2019) introduced CLUTRR, a benchmark assessing LLMs’ structural learning by inferring kinship relations in stories, requiring relationship extraction and logical rule inference. Zhu et al. (2024) proposed the Hypotheses-to-Theories (HtT) framework to improve LLM reasoning by learning explicit rules in two stages: generating and verifying rules (induction) and using the obtained rule library for reasoning (deduction). HtT enhances relational and numerical reasoning and concept learning.
Madaan et al. (2023) introduces a novel technique for improving machine learning models through iterative refinement. This approach allows models to improve their performance by continuously evaluating and adjusting their predictions based on self-generated feedback. By critiquing their own outputs, models can identify errors and make corrections over successive iterations, leading to increased accuracy and robustness across different tasks. Our work builds upon this approach by employing a formal method to evaluate and then refine itself.
In a related study, Dziri et al. (2023), the authors investigate the limitations of transformer models in handling composition tasks. Their results show that, despite their strengths, transformers face significant challenges in dealing with compositionality, which involves understanding and generating complex structures from simpler components. This limitation highlights the need for innovative approaches, such as self-refining, to further enhance the capabilities of machine learning models.

### v2

# Inductive Learning of Logical Theories with LLMs: An Expressivity-Graded Analysis

@@ Abstract
This work presents a novel systematic methodology to analyse the capabilities and limitations of Large Language Models (LLMs) with feedback from a formal inference engine, on logic theory induction. The analysis is complexity-graded w.r.t. rule dependency structure, allowing quantification of specific inference challenges on LLM performance. Integrating LLMs with formal methods is a promising frontier in the Natural Language Processing field, as an important avenue for improving model inference control and explainability. In particular, inductive learning over complex sets of facts and rules, poses unique challenges for current autoregressive models, as they lack explicit symbolic grounding. While they can be complemented by formal systems, the properties delivered by LLMs regarding inductive learning, are not well understood and quantified. Empirical results indicate that the largest LLMs can achieve competitive results against a SOTA Inductive Logic Programming (ILP) system baseline, but also that tracking long predicate relationship chains is a more difficult obstacle than theory complexity for LLMs.
Introduction
The integration of Large Language Models (LLMs) with formal methods stands out as a promising frontier in the field of Natural Language Processing. It is an avenue for improving model inference control and explainability, by both complementing the content flexibility of Large Language Models (LLMs) with the systematicity of symbolic/formal systems (Quan et al. 2024a, b) and by using well-defined formal settings to assess the underlying inference properties of the model.
Inductive Logic Programming (ILP) is a subfield of symbolic AI which focuses on methods that can derive (generalise) theories to explain observed facts and rules (Muggleton 1991; Nienhuys-Cheng and de Wolf 1997). Addressing inductive learning over complex sets of facts and rules, poses unique challenges for current autoregressive LLMs, as they do not operate over data symbolically, rather combining an extensive set of structural an semantic signals to approximate the most probable answer in a generative fashion.
@@ Introduction
RQ1. To what extent the combination of an LLM with feedback from a formal inference engine can compare to a SOTA inductive logic programming (ILP) system in logic theory induction, w.r.t. inference quality at different expressivity levels?
RQ2. How does the expressivity of the target theories affect inference quality of LLMs for inductive reasoning?
Figure 1: The proposed method to evaluate theory induction with an LLM in Prolog based on background knowledge and training examples. The process starts with a prompt generator $(c)$ that formulates prompts for an LLM $(a)$. Both the background knowledge and training sets are parameterised by different noise and rule expressive power levels: Chain, Rooted Directed Graph (DG), Disjunctive Rooted DG, and Mixed. The LLM generates theories, which are then evaluated by a logic program interpreter $(b)$. The evaluation feedback, including accuracy, precision, recall, and F1 scores, as well as wrongly classified examples, is used to refine the prompts iteratively. We analyse and categorise the generated theories according to their expressive power $(d)$.
In order to address these RQs, we propose a method for combining iterative theory refinement on stock LLMs (i.e., zero-shot inference (Brown et al. 2020; Radford et al. 2019)), a formal ILP inference engine and a synthetic generator for inductive reasoning datasets, in order to perform systematic evaluations of the LLM induced theories, using state-of-the-art ILP solvers as a baseline. Moreover, to quantify the extent in which LLMs can address ILP tasks, the evaluation is graded w.r.t. the expressivity of the target rulesets. Figure 1 schematises the proposed approach.
@@
    (Empirical) A detailed empirical analysis on the strengths and limitations of SOTA LLMs regarding logic theory induction generation, according to target ruleset expressivity.
@@
The remainder of the paper is organised as follows: Section “Inductive Learning, Expressive Power & Datasets” formalises the relevant ILP concepts, dataset generation, and LLM used in the paper. Section “Proposed Approach” describes the proposed method and algorithms. Section “Experiments” presents the experimental procedures, results and discussion, followed by related work and conclusions.
Inductive Learning, Expressive Power & Datasets
In this section we introduce the target task and the typology of inductive learning expressivity levels, as well as the dataset generation process.
Inductive Logic Programming
@@
Given:
  • Background Knowledge (BK): A set of logical clauses that represent prior knowledge about the domain.
  • Positive Examples ($E^{+}$): A set of ground facts (instances) which the learned theory should entail (i.e., these are examples that should be true according to the theory).
  • Negative Examples ($E^{-}$): A set of ground facts which the learned theory should not entail (i.e., these are examples that should be false according to the theory).
Goal: Find a hypothesis $H$ (a set of logical clauses) such that:
  1. 1.
    Completeness: For every example $e\in E^{+}$:
    $H\cup BK\models e$
    Meaning the hypothesis $H$, together with the background knowledge $BK$, should entail all positive examples.
  2. 2.
    Consistency: For every example $e\in E^{-}$:
    $H\cup BK\not\models e$
    Meaning the hypothesis $H$, together with the background knowledge $BK$, should not entail any negative examples.
Formally: An ILP system seeks a hypothesis $H$ that satisfies:
@@
Conclusion: The learned hypothesis $H$ should thus be a logical theory that explains the positive examples and excludes the negative examples, based on the given background knowledge.
Theory Expressivity
Inductive learning can be organised according to different classes of expressive power, which involves a typology of the structural complexity of the problem. Moreover, variables such as the amount and type of noise within the evidence set (such as the number of incorrect or missing facts or completeness levels) can be integrated within this setting. Following the typology of (Cornelio and Thost 2021) four base categories of rules can be introduced based on their dependencies: Chain, Rooted Directed Graph (RDG), Disjunctive Rooted DG, and Mixed. Each category represents a hierarchical generalisation, with each step encompassing a broader range of structures. Starting with CHAIN, which represents a linear composition of predicates, RDG is a generalisation of CHAIN, meaning it includes all chain structures but also accommodates more complex dependencies. Moving up, DRDG generalises RDG by incorporating directed relationships, thus offering a more extensive representation of dependencies. Finally, Mixed is a DRDG with recursion, containing connected components from CHAIN, RDG, and DRDG. Each progression from CHAIN to MIXED represents a step towards greater inclusivity and complexity in the types of structures captured.
The expressive power of each class and their characteristics are summarised in Table 1. A detailed description of each class is provided in the supplementary material [^1] A version of this paper with the supplementary material can be found at https://arxiv.org/pdf/2408.16779. .
@@
Table 1: Characteristics for each dataset category. #Parents refers to the number of times a predicate appears as the head of a rule in the context of rule learning and logical induction. It indicates the number of rules through which each rule can deduce relevant facts. Recursive refers to whether a predicate in the head of a rule can also occur in the body of the same rule. Alt. rules indicate whether a predicate can be deduced by alternative rules.
Dataset synthesis
In order to generate datasets for rigorous analysis, this study employed the RuDaS tool (Cornelio and Thost 2021) to systematically vary parameters such as noise, open-world degree, and missing data. By adjusting these factors in conjunction with the category parameter, it is possible to ensure comprehensive coverage of different structural configurations and expressivity levels.
@@
  • The minimum and maximum number of Directed Acyclic Graphs (DAGs), which refer to the range of distinct rule structures generated within each dataset, were both set to 1 (mindags = 1, maxdags = 1).
  • Noise levels, which include the addition of irrelevant facts or the removal of relevant ones that do not support the rules, were systematically varied at intervals of 0.1, 0.2, and 0.3.
  • The percentage of missing data (missing), which determines the proportion of data entries - specifically consequences - that are intentionally left out, and open world degree (owa) were similarly varied across 0.1, 0.2, and 0.3.
  • The category parameter was set to cover all classes of expressivity described in the previous section, and listed in Table 1. The distribution of each category in the synthesised dataset is an independent hyperparameter, discussed in the experiments section.
[…]
Proposed Approach
The proposed approach can be divided in two parts: iterative refinement and graded evaluation. They form a systematic evaluation loop, covering all the expressivity classes described in the previous section, for a given set of LLMs and dataset synthesis parameters.
Iterative Refinement
@@
(a) A language model capable of generating a theory $H$, based on background knowledge, positive and negative examples, and hypothesis search, provided as a prompt text in a logic program language. Typically an LLM with structured (e.g., code) generation capabilities. (b) A logic program interpreter. We use Prolog (Warren et al. 2023) as the logic program language. (c) A prompt generation component, that interleaves logical programming language with natural language queries designed to drive the theory induction responses. The logical programming language expresses background knowledge and the relevant outputs of the program interpreter. (d) An evaluation module, that uses the logic program interpreter to execute the generated theory $H$ as logical rules, and computes a set of evaluation metrics.
@@ 3. Evaluation of Hypothesis: Checking for the completeness and consistency conditions:
from which accuracy (ACC) ($\frac{TP+TN}{TP+TN+FP+FN}$), precision (P) ($\frac{TP}{TP+FP}$), recall (R) ($\text{Recall}=\frac{TP}{TP+FN}$) and F1-score (F1) ($2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$) can be generated.
4. Theory refinement: Following an initial evaluation, the LM is tasked to refine the induced theory iteratively. Each refinement round involves adjusting the theory based on feedback from the Prolog interpreter validation. The refinement aims to improve the theory’s performance by addressing misclassifications and enhancing its predictive capabilities. If $H$ does not satisfy completeness and consistency, update the input representation based on feedback and generate a new hypothesis using the language model, given the Feedback Context $\leftarrow\{FP,FN,P,R,F1,ACC\}$ and the final prompt input Input $\leftarrow$ (Refinement Prompt + Context + Feedback Context): $H\leftarrow\mathcal{LM}(\text{Input})$.
@@
Graded evaluation
A synthetic data generator is used to control for the input parameters of the ruleset expressive power, namely: categorical distribution (CHAIN, RDG, DRDG, etc.), background knowledge, positive examples and negative examples as well as the amount of noise introduced within the dataset.
1. Categorised Learning Sets: Consisting of $C$: set of ruleset expressivity classes (e.g., CHAIN, RDG, DRDG, etc.), $N$: set of noise levels, $S$: number of samples per combination of $C$ and $N$. For each $c\in C$ and $n\in N$, generate $S$ datasets $\{D_{c,n,i}\mid i=1,\ldots,S\}$ where each dataset $D_{c,n,i}$ includes:
@@
Experiments
In order to answer the research questions, a set of experiments was elaborated to systematically analyse the theory inducing capabilities of a set of most popular open-source LLMs and two versions of GPT, with the proposed approach, having the state-of-the-art ILP system Popper (Cropper and Morel 2021) as a baseline. The tests covered all data categories discussed on the section “Inductive Learning, Expressive Power & Datasets”, allowing a graded analysis w.r.t. the expected level of induction expressivity and tolerated noise.
Experimental Setup & Dataset
For each data category, five datasets were generated using RuDaS (Cornelio and Thost 2021). The size of each dataset was set to $XS$ ($min=50$, $max=100$, $support=3$), and noise, missing, and open world were all set to 0.1, then all set to 0.2, and finally all set to 0.3. This resulted in 105 datasets in total, with 35 datasets for each rate. Subsequently, two methods are used to induce a theory for each dataset: (1) employing Popper, with NuWLS (Chu, Cai, and Luo 2023) and WMaxCDCL, varying its time limit parameter from 10 to 800 seconds; (2) applying the proposed iterative LM theory refinement method (Section “Proposed Approach”), with parameters $Max_{iter}=4$ and $MT_{thresh}=1.0$. Three different LLM models were used for (2): Open AI [^2] https://openai.com/’s model GPT-4o [^3] https://openai.com/index/hello-gpt-4o, Mistral AI [^4] https://mistral.ai/’s Mixtral-8x7B [^5] https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 (Jiang et al. 2023), and Google’s Gemma [^6] https://huggingface.co/google/gemma-7b-it (Team et al. 2023).
@@
Table 2: Statistics for each dataset category. A detailed description of each can be found in Section “Inductive Learning, Expressive Power & Datasets”. The mean $\pm$ standard deviation (mean $\pm$ std) for each category is provided
@@
[…]
Results & Discussion
[…]
Figure 2: F1 score trends across categories. Different models (GPT-4o, Llama3 8b instruct, Popper, and Mixtral-8x7B-Instruct-v0.1) under varying noise levels and categories reveal distinct performance patterns. GPT-4o demonstrates stable accuracy yet sensitivity to noise, particularly in complex rule-based categories like RDG and DRDG. Mixtral-8x7B-Instruct-v0.1 exhibits mixed results with notable variability across categories particularly in more complex tasks. Llama3 8b instruct delivers lower scores, indicating challenges in reasoning and theory generation.
@@
Inducing theories on long relation chains is the major obstacle for LLMs, rather than the expressivity level.
With the CHAIN category being the least expressive and one of the most consistently solved by Popper, but none of the tested LLMs was able to overcome the baseline performance at all noise levels (Figure 2 CHAIN category). This suggests that such models have a limited capacity of tracking long relationship chains of independent predicates. This addresses a part of the answer for RQ2.
@@
Upon increasing the iteration limits from 1 to 4, it was found that the metrics can either increase or decrease non-monotonically. No significant overall performance improvements were observed beyond 3 iterations. Thus $Max_{iter}$ was set to 4 and the iteration with the best accuracy is taken as the final result.
Performance is remarkably lower on more expressive rule sets at moderate noise levels.
[…]
@@
While Popper consists of a combination of serial algorithms, the parallel nature of transformer-based LLMs allows them to operate at times about 3 orders of magnitude lower, given enough computational resources. This can be observed in Figure 3, for all classes and all tested noise intervals.
Error analysis
@@
prolog theory :- p(X, Y), pos(p0(X, Y)) - positive. p(X, Y), neg(p0(X, Y)) - negative. \+ p(X, Y), pos(p0(X, Y)) - false. \+ p(X, Y), neg(p0(X, Y)) - true.
@@
[…]
@@
Related Work
Neural symbolic computation combines neural networks with symbolic methods to enhance AI reasoning capabilities. Yang and Cohen (2017) introduced Neural Logic Programming, An end-to-end differentiable model integrating neural networks with logic programming. Within the LLM-Symbolic space, Wan et al. (2024) developed LogicAsker, which evaluates and improves LLMs’ logical reasoning using propositional and predicate logic. It identifies reasoning failures and enhances capabilities through in-context learning. Within the context of symbolic toolformers over LLMs, Quan et al. (2024a; 2024b) proposed methods of improving explanatory reasoning with the support of formal iterative cycles using both logical solvers and theorem provers for supporting more controlled step-wise reasoning.
Despite these advancements at the interface of LLM-based reasoning and formal controls, it is unclear the extent and the conditions in which LLMs can perform formal reasoning (Huang and Chang 2023). Sinha et al. (2019) introduced CLUTRR, a benchmark assessing LLMs’ structural learning by inferring kinship relations in stories, requiring relationship extraction and logical rule inference. Zhu et al. (2024) proposed the Hypotheses-to-Theories (HtT) framework to improve LLM reasoning by learning explicit rules in two stages: generating and verifying rules (induction) and using the obtained rule library for reasoning (deduction). HtT enhances relational and numerical reasoning and concept learning.
Madaan et al. (2023) introduces a novel technique for improving machine learning models through iterative refinement. This approach allows models to improve their performance by continuously evaluating and adjusting their predictions based on self-generated feedback. By critiquing their own outputs, models can identify errors and make corrections over successive iterations, leading to increased accuracy and robustness across different tasks. Our work builds upon this approach by employing an ILP solver to evaluate and then using its output as critique for the refinement process.
In a related study (Dziri et al. 2023), the authors investigate the limitations of transformer models in handling composition tasks. Their results show that, despite their strengths, transformers face significant challenges in dealing with compositionality, which involves understanding and generating complex structures from simpler components. This limitation highlights the need for innovative approaches, such as self-refining, to further enhance the capabilities of machine learning models.